OpenStudy (anonymous):

Differentiate h(theta)=5 theta sec theta

OpenStudy (amistre64):

its a product rule ... h = 5x sec(x)

OpenStudy (anonymous):

so would that sec(x) become a sec(x)tan(x)

OpenStudy (amistre64):

yes

OpenStudy (anonymous):

swell.

OpenStudy (amistre64):

h = 5x sec(x) h' = 5x' sec(x) + 5x sec(x)' h' = 5sec(x) + 5x sec(x)tan(x) h' = 5sec(x) (1+x tan(x)) or however you want to present it :)

OpenStudy (anonymous):

thanks a bunch!

OpenStudy (anonymous):

How did you get the one in "(1+x tan(x))"?

OpenStudy (amistre64):

i factored out a common factor ...

OpenStudy (amistre64):

5sec(x) + 5x sec(x)tan(x) let a = 5 sec(x) a + a xtan(x) factor out the a a (1+ x tan(x))

OpenStudy (anonymous):

I think I'm understanding that now...thanks again! :)